### v3

# Generalizing Medical Image Representations via Quaternion Wavelet Networks

Abstract
Neural network generalizability is becoming a broad research field due to the increasing availability of datasets from different sources and for various tasks. This issue is even wider when processing medical data, where a lack of methodological standards causes large variations being provided by different imaging centers or acquired with various devices and cofactors. To overcome these limitations, we introduce a novel, generalizable, data- and task-agnostic framework able to extract salient features from medical images. The proposed quaternion wavelet network (QUAVE) can be easily integrated with any pre-existing medical image analysis or synthesis task, and it can be involved with real, quaternion, or hypercomplex-valued models, generalizing their adoption to single-channel data. QUAVE first extracts different sub-bands through the quaternion wavelet transform, resulting in both low-frequency/approximation bands and high-frequency/fine-grained features. Then, it weighs the most representative set of sub-bands to be involved as input to any other neural model for image processing, replacing standard data samples. We conduct an extensive experimental evaluation comprising different datasets, diverse image analysis, and synthesis tasks including reconstruction, segmentation, and modality translation. We also evaluate QUAVE in combination with both real and quaternion-valued models. Results demonstrate the effectiveness and the generalizability of the proposed framework that improves network performance while being flexible to be adopted in manifold scenarios and robust to domain shifts. The full code is available at: https://github.com/ispamm/QWT.
Index Terms:
Generalizable Neural Networks Quaternion Wavelet Transform Task-Agnostic Deep Learning Quaternion Neural Networks Medical ImagingI Introduction
The generalizability challenge in neural network deployment is still a wide-open problem, being amplified by the increasing number of available image datasets coming from different sources and diverse acquisition processes. This matter becomes crucial when dealing with medical image datasets, whose samples are generated by a variety of hospitals, devices, and operators, thus similar datasets can indeed have wide variations. Data of the same medical exam obtained in diverse hospitals may suffer from a domain shifting affecting the generalizability of pretrained models that often have to be re-trained from scratch on the new data [1]. In fact, deep learning techniques have profoundly permeated the field of medical image analysis [2, 3, 4], reaching impressive results in manifold applications, including detection and classification [5, 6, 7], segmentation [8, 9], reconstruction, and inverse problems solution [10, 11, 12], among others. However, such approaches are usually task- or data-specific, thus losing generalizability and undermining their adoption in different contexts. Furthermore, medical images are usually high-quality and full-of-details images that have to be resized and downscaled to be handled by neural networks due to computational limits. This process may compromise image quality and cause a loss of crucial details even in the region of interest, which is often a tiny portion of the whole picture. Due to such image daintiness, the preprocessing often plays a pivotal role in medical analysis [13], being a mandatory step before developing any deep learning model for image datasets [14].
As part of preprocessing, techniques for identifying important and explanatory training examples have been developed for natural images [15, 16] aiming at discarding superfluous data samples and improving the generalization ability of neural networks. Similarly, several works try to establish the most representative patches (or whole sample) to be involved in the training of medical deep learning approaches showing the ability to improve models performance [17, 18]. Simultaneously, a further approach consists of fusing different image modalities in a single and more informative sample [19, 20]. This method is particularly suitable when the dataset contains multimodalities or multiresolution samples. Fusing different modalities combines important information coming from the various acquisitions and may build a more detailed structure, improving pathology examination accuracy. A common approach for medical image fusion aims at leveraging wavelet transforms that split the original image into different frequency components, providing different levels of detail [21, 22]. Recently, a further improvement has been brought by the adoption of the quaternion wavelet transform (QWT) that provides a hypercomplex approach with additional sub-bands and several advantages, including shift invariance, over the standard real-valued wavelet [23, 24, 25, 26].
In this paper, we introduce a novel generalizable framework for medical image analysis and synthesis, referred to as quaternion wavelet network (QUAVE). The proposed approach is data-agnostic, and it does not suffer from domain shifting in data, thus it can be employed with any kind of medical image dataset, for any small and large-scale neural model, and for different tasks. We integrate a quaternion-based preprocessing in a real-valued analysis procedure, to leverage quaternion capabilities in any real-valued scenario, while simultaneously further generalizing the adoption of quaternion and hypercomplex neural networks. QUAVE is built upon two core blocks. The first block exploits quaternion wavelet potentials to extract salient features from medical images that are usually full of details and require detailed preprocessing to preserve crucial content characteristics. These features are then processed by the second QUAVE block, where a neural model extracts the most informative content from the QWT bands to build enhanced representations that can be involved as input to any model for image analysis. Indeed, thanks to the QWT low-frequency and high-frequency sub-bands, we make any model for analysis aware of both approximations and details of the input image, improving its ability to reconstruct or generate images full of details, while also enhancing segmentation capabilities. The overall structure of our framework is depicted in Fig. 1. QUAVE’s output can be generalized to other datasets without requiring the repetition of the training procedure. Furthermore, since the proposed wavelet-based preprocessing is performed in the quaternion domain, we can easily seize quaternion-valued neural networks (QNNs) for the learning stage to further exploit the capabilities of hypercomplex algebra. Indeed, QNNs have been proven to achieve interesting results in processing natural images [27, 28, 29, 30, 31]. Due to the four-dimensional nature of quaternions, QNNs properly handle inputs with dimensions/channels, thus restricting the use of QNNs to a few sets of data. Instead, through QUAVE, we generalize the usage of quaternion and hypercomplex neural networks for single-channel inputs while also improving performance. In fact, by exploiting the proposed approach it is possible to easily handle grayscale and one-channel images in QNNs. Indeed, through the QUAVE processing, we obtain four sub-bands that can be easily encapsulated in a quaternion. Due to the obvious correlations among these sub-bands, QNNs equipped with our method can improve performance by grasping such relations and sharing information among channels.
We prove our theoretical claims through a wide experimental evaluation covering three different tasks and three common medical image datasets. In every test we conduct, QUAVE enhances model performance outperforming previous techniques, according to a plethora of objective metrics and a visual inspection. Therefore, the proposed approach is flexible enough to be adopted with any kind of dataset, especially if it contains images full of details that can be exploited by high-frequency sub-bands.
The rest of the paper is organized as follows. Section II introduces basic concepts of quaternion algebra, Section III expounds on the wavelet and quaternion wavelet transforms for images, and Section IV presents the proposed selection method. The experimental setup is then introduced in Section V and validated in Section VI. Finally, conclusions are drawn in Section VII.
II Fundamentals of Quaternion Learning
In this section, we expound on basic concepts of quaternion algebra, the foundations of quaternion neural networks and how they can be involved in a variety of studies. Indeed, real-valued data can be easily processed with quaternion-valued models by encapsulating input dimensions into the components of a quaternion. Similarly, a quaternion framework can be handily integrated into a real-valued analysis.
II-A Quaternion Algebra
The quaternion set of numbers belongs to the class of Clifford algebras, being a four-dimensional associative, normed, and non-commutative division algebra over real numbers. While complex numbers are defined by a real-valued component and one imaginary unit, quaternion numbers comprise two additional imaginary components, being defined as:
whereby are real-valued coefficients and imaginary units, which comply with the properties:
1.
2.
3.
where is the vector product in . An alternative quaternion representation can be derived considering the magnitude and the three phases:
in which is the amplitude of the quaternion and are the phase angles of it. The latter can be computed as follows:
The quaternion domain is endowed with a norm , a conjugate , the operations of scalar multiplication , and associative multiplication of elements.
Due to the vector product non-commutative property, the Hamilton product has been introduced for properly modelling imaginary unit interplays [32]:
II-B Learning in the Quaternion Domain
Quaternion neural networks (QNNs) are built upon the Hamilton product for multiplying and convolving weights with input data [33, 34]. More precisely, a real-valued fully connected layer with weights matrix and bias has the following expression:
where the real-valued input is modelled into the real-valued output . A quaternion fully connected (QFC) layer involves quaternion characters and its definition is based on the Hamilton product between the weights matrix, arranged in the form of (4), and the quaternion input as:
The QNNs formulation through the Hamilton product has two key advantages. First, since the quaternion weight matrix is composed of four sub-matrices , with , each with the parameters of the complete matrix , and these sub-matrices are reused to build the final weight matrix according to (4), QNNs save of free parameters with respect to real-valued counterparts. Second, due to such sharing of weight sub-matrices, each parameter is multiplied by each dimension of the input (e.g., by each channel of RGB images, or of multichannel signals), thus capturing complex relationships among input dimensions and preserving their correlations [35, 36]. This allows QNNs to gain comparable results when processing multidimensional data despite the lower number of free parameters. However, due to the four-dimensional nature of quaternion numbers, QNNs are mainly limited to multidimensional signals, while their extension to one-channel and monodimensional data is still an open problem.
III Quaternion Wavelet Representation of Medical Images
In this section, we expound on the proposed approaches for extracting salient characteristics from input images. Moreover, we also explain how to leverage these features with quaternion neural networks.
III-A Quaternion Embedding of Real-Valued Wavelet Representations
The 1D discrete wavelet transform (1D-DWT) characterizes a signal by shifting a scaling function and shifting and scaling a wavelet function [37]. In case of 2D signals, the 2D-DWT can be obtained by involving tensor products of 1D-DWTs over the two dimensions, resulting in the scaling function and three wavelets , , and that extract diagonal, horizontal and vertical features, respectively [38, 39]. Through this decomposition, the 2D-DWT (from now on, DWT for simplicity) of an image allows for the derivation of a low-frequency (LL) sub-band and three high-frequency (LH, HL, HH) sub-bands. However, this kind of decomposition has some disadvantages. First of all, it lacks translation invariance. Secondly, it does not contain image phase information that usually cares about describing the spatial information of image contents. Therefore, even small shifts in the image content, which are quite common in datasets acquired through different devices or by diverse operators, can crucially affect wavelet magnitude [40, 41]. Indeed, phase information may be crucial in medical images that could contain artifacts due to patient movements or inaccurate equipment. Consequently, in order to build a more powerful representation of DWT image features, we propose to leverage quaternion algebra properties by encapsulating the four sub-bands in a quaternion as and then exploiting its magnitude and phases characterization in (2). The phase angles can be then computed following (3). In this way, the input image is first decomposed into different frequency sub-bands by means of a DWT and then processed through quaternion algebra to get also the missing phase information, building a more accurate representation. Such manipulated data can be then fed into real, quaternion, or hypercomplex-valued neural networks without needing additional preprocessing.
III-B Enhanced Representation via Quaternion Wavelet Transform
An enhanced representation of medical images, that we embed in our QUAVE framework, involves the use of the dual-tree quaternion wavelet transform (QWT) [39]. It comprises four quaternion wavelet transforms that are built by means of a real DWT and its three Hilbert transforms on the and axis. The QWT decomposes the input image in four quaternion wavelet sub-bands for a total of 16 real sub-bands. As for the DWT, the transformation result comprises a scaling function and three wavelets in the vertical, horizontal and diagonal direction, respectively [42, 41]. More formally, they are defined by:
whereby are a real-valued filter and its corresponding Hilbert transform.
More concretely, the QWT of an image can be computed by combining four filters. To this end, an asymmetric, orthogonal and biorthogonal Daubechies wavelet with 8 vanishing moments (db8) can be considered, following [43]. From the db8, the decomposition low-pass () and high-pass filters ( are extracted, and then the Hilbert transform is applied to them to obtain the filters and . A plot of such filters is shown in Fig. 2. Once and are available, the QWT is calculated by combining filters and downsampling operations according to Fig. 3. The result comprises low-frequency and approximation sub-bands () and high-frequency sub-bands (). Furthermore, the four components in (7) can be organized into a quaternion, whose square magnitude is non-oscillatory, meaning that the transform is approximately shift-invariant [39].
Among these featured images, some previous works suggest only involving the four low-frequency components as input to the model since, under an empirical evaluation, they improve classification accuracy [44, 45]. However, these tests were conducted with vanilla classifiers without involving deep learning approaches that are notoriously more powerful. Indeed, shallow models may not have enough capacity to handle very complex inputs such as high-frequency wavelet details. Neural networks, instead, are more complex systems and more detailed inputs can further enhance model expressiveness, improving performance. Furthermore, for the task of image classification, there may be sufficient to have an approximate representation of the input, while a lack of high-frequency details may undermine results on other tasks such as reconstruction, segmentation or generation. On the other hand, high-frequency features are less informative with respect to low ones, thus a proper combination of low and high sub-bands can better represent input images. However, involving the sixteen sub-bands as input to a neural model may undermine its performance due to redundant features, forcing the model to distinguish between relevant and uninformative channels before effectively learning a proper representation. Furthermore, QNNs require four-channel inputs to be encapsulated into a quaternion, thus it is not suitable to directly involve the sixteen sub-bands.
To this end, the QUAVE framework involves a selection strategy, detailed in the next section, to choose the most explanatory features to be fed as input to real and quaternion-valued networks.
IV The Proposed QUAVE
In this section, we expound upon our pathway for extracting the most informative representation from the sixteen features obtained through the Quaternion Wavelet Transform (QWT). While the Discrete Wavelet Transform (DWT) produces four sub-bands that can be directly utilized as input for models, the QWT generates sixteen sub-bands, potentially resulting in redundant information if they are used as input to the network. Moreover, quaternion-valued models require exactly four-channel inputs, rendering the direct conveyance of the sixteen QWT sub-bands to the network impractical. To address these challenges, we propose a learning-based approach to build the most informative representation among the sixteen quaternion wavelet sub-bands for involving it as enhanced input to neural models. Recent literature demonstrates the utility of wavelets in various tasks, such as image fusion methods [25] involving multimodal image fusion using wavelets and multiple features. Another area of research utilizes wavelets to enhance super-resolution by enriching output images with improved fine-grain details [46]. Among this class of models, the super-resolution approach comprises a series of image enhancement techniques, including feature refinement [47], which hold significant value for our intended purpose. Based on this premise, we propose a neural model that leverages quaternion wavelets to perform super-resolution, extracting the most informative content by means of the quaternion wavelet features. The underlying idea involves employing transfer learning, wherein QUAVE is trained to perform a super-resolution task on a generic medical dataset (e.g., IXI). Subsequently, the pretrained QUAVE model is utilized to extract the features for a downstream task of interest, such as reconstruction, segmentation, or image translation. Extracting the features entails employing the pretrained QUAVE model and, as depicted in Fig. 4, selecting the weights from the second-last layer, prior to the inverse wavelet transform. The architecture of QUAVE consists of two principal modules. The rough feature extraction module (RFEM) provides the necessary features for performing the quaternion wavelet transform. Subsequently, the wavelet transform module takes the features extracted by the previous modules, it first applies the QWT. Moving forward, the module operates in four groups: one low-frequency group and three high-frequency groups. Before entering the feature refinement block, we reduce the dimensionality of the image features by a factor of four using a convolutional layer. To extract the enriched features of the sub-bands, two linear layers are inserted before the inverse quaternion wavelet transform, which outputs the final super-resolution image. Figure 4 presents an overview of the proposed architecture, which comprises multiple residual groups [48]. Each group consists of residual blocks with convolutional layers and ReLU activation functions. Our methodology allows us to retrieve the features, representing four sub-bands of the quaternion wavelet, that are considered the most informative by the neural model, as they activate the neurons that yield the best performance. During the training phase of a task of interest (e.g. segmentation), we perform inference on pretrained QUAVE, since we utilize the four sub-bands that the model outputs as input for the task of interest. The features representing four sub-bands of the quaternion wavelet can naturally represent a quaternion number, and this fits perfectly for our analysis of this methodology on quaternion-valued neural networks.
V Experimental Setup
We perform experiments in multiple tasks and with different benchmark datasets to show the generalizability of our framework and to strengthen our theoretical claims. To evaluate the effectiveness of our approach, we utilize QUAVE pretrained on a medical grayscale image dataset. We first consider tasks regarding image analysis such as reconstruction and segmentation and then an image synthesis task, that is image modality translation with a large-scale model. We conduct analysis on two algebraic domains: first, we set up experiments to prove that our framework can enhance the performance of real-valued neural networks. Second and most importantly, we perform evaluations with quaternion-valued models to show that QUAVE outperforms other preprocessing techniques and it is flexible to be adopted in different domains, further improving QNNs generalizability. Moreover, we consider diverse types of input images. We take into account medical grayscale and RGB images, and multimodal images, proving that the proposed method can be generalized to any kind of input, thus being data-agnostic and not suffering from domain shifting.
V-A Description of the Tasks
V-A1 Reconstruction
As a first step of our evaluation, we take into account the IXI dataset111https://brain-development.org/ixi-dataset for the reconstruction task. It comprises magnetic resonance images collected from three different hospitals in London, and it is organized as follows: T1, T2, PD-weighted images, MRA images and diffusion-weighted images. We consider the combination of the multimodal T1 and T2 images, namely brain Nuclear Magnetic Resonances (NMR) of and patients respectively. These are provided in NIFTI format, which is a typical format of neuroimaging. In particular, each 3D NRM is composed of different images, each of which identifies a portion of the brain. The sequence of these images is a temporal scan of the area analyzed, so the total number of images in each 3D NRM varies according to the acquisition. For example, in T1 it goes from to , while in T2, it is from to . We preprocess images by keeping only one significant image for the patient into account. After this step, we obtain images from the combination of T1 and T2 ones, each of size . We employ of them for training and for validation.
V-A2 Segmentation
The second set of experiments comprises the segmentation task on the Kvasir-SEG dataset [49]. This dataset contains snapshots of polyps with manually annotated masks verified by an experienced gastroenterologist. The samples range from to pixels, and we resize them to . We involve samples for training and the remaining part for validation. Interestingly, and differently from previous datasets, the Kvasir-SEG dataset contains RGB images with three channels. Therefore, we train the real-valued model with original three-channel samples, while we apply the DWT and the QWT computed on the corresponding grayscaled image.
V-A3 Image Modality Translation
The third set of experiments concerns image modality translation on the CHAOS - Combined (CT-MR) Healthy Abdominal Organ Segmentation dataset. It contains CT and MR scans from unpaired abdominal image series, the ground truth is generated with the auxiliary participation of multiple experts’ annotations and introduced for the first time in [50]. The data comes from patients from the Department of Radiology of Dokuz in Turkey, half of them have done only a single CT scan while the others went through MR scans. For our experiments we have used all the data resizing the slices at , obtaining approximately samples, we randomly select from all the modalities to be the training dataset, and the remaining as test one.
V-B QUAVE in Combination with Real- and Quaternion-Valued Models
We involve different models in our tests. For the reconstruction task, we consider a vanilla convolutional autoencoder (CAE) [35], whereby the encoder is composed of three convolutional layers with kernels for the IXI dataset. The decoder has a mirrored structure with transposed convolutions. The quaternion CAE (QCAE) [35] has the same backbone while involving quaternion convolutions instead of real-valued ones. For each reconstruction experiment, we set the learning rate of the Adam optimizer to , the batch size to , and the number of epochs to , repeating the learning process for multiple runs with different seeds.
Concerning the segmentation task, we consider a U-Net [51] and a more advanced U-Net++ [52] in our experiments, as they are widespread models for this scenario. The U-Net encoder comprises four blocks, each with two convolutions, batch normalization, and ReLU activation function, while max pooling is interleaved between each block. The number of filters increases going deeper, being , each with kernel sizes and stride . The bottleneck adds a further block with filters. Then, the information is passed to the decoder, which has the same architecture as the encoder except for the replacement of transposed convolutions instead of max pooling operations. A final convolutional layer refines the resulting segmentation map with channel. Similar to the QCAE, the quaternion U-Net (QU-Net) replaces convolutions and transposed convolutions with quaternion-valued counterparts. The more advanced U-Net++ has a similar structure to the original one, inserting however dense connections that aggregate intermediate feature maps. Every hyperparameter is set according to the paper [52]. In these experiments, as for the reconstruction task, we set the learning rate to for the U-Net and to for the U-Net++, as suggested in the paper [52], with Adam optimizer and the batch size to . We train all the models for epochs, repeating the learning process for multiple runs with different seeds.
Regarding the image modality translation task, we consider the recent Target-Aware Generative Adversarial Network (TarGAN) [12]. The TarGAN translates images from a source domain to a target domain while paying particular attention to the region of interest passed as a segmentation map during the translation. The generator takes as input the original image and the four wavelets associated with it, the target area of interest, and the output target modality, generating in the output the translated image in the desired domain. This network comprises two branches, the first one for the whole image and the second one for the region of interest. A single branch is composed of an initial encoder, then a set of blocks shared with the other branch, and finally, a decoder to output the desired image. This network stacks convolutional blocks that employ a convolution layer then the instance normalization, and finally a Leaky ReLU activation function. The upsampling blocks instead involve upsampling layers by interpolation suddenly a convolution operation, and, as in the previous block, finally the instance normalization with Leaky ReLU. The discriminator has a similar structure, for each detail regarding the model’s structure and hyperparameters we refer to the original paper [12]. Overall, this model comprises million parameters, being the largest in our experiments. In this case, we introduce the quaternion-valued counterparts as the quaternion TarGAN has not been introduced yet in literature.
V-C Metrics
As objective evaluation metrics in the reconstruction task, we compute the structural similarity index (SSIM), the mean squared error (MSE), and the Fréchet inception distance (FID). While SSIM measures image degradation through structural information, FID estimates how the distributions of real and reconstructed images are far from each other.
For evaluating the results of the segmentation task instead, we compute the Dice similarity coefficient (DSC), the mean intersection over union (mIoU), and the mean absolute error (MAE). In order to perform a robust evaluation of the image modality translation task, we employ both a standard FID and an FID proposed in the original work [12]. The latter is a composition of images (16 slices) passed into a different model [53] and then evaluated with the Fréchet distance. We carefully follow the recipe in [12] for the computation of this task-specific metric.
VI Results and Discussion
In this section, we present the selected quaternion sub-bands we employ in our tests and the empirical evaluation we accomplish for validating our theoretical claims. We perform multiple runs with different seeds for each experiment and we report in the tables the average score and the standard deviation for each metric.
| Model | Params | SSIM | MSE | FID |
| CAE[35] | 0.9915 0.007 | 0.0006 0.0001 | 0.0046 0.0035 | |
| QUAVE-CAE | 0.9915 0.005 | 0.0006 0.0001 | 0.0036 0.0006 | |
| QCAE[35] | 0.9504 0.002 | 0.0015 0.0001 | 0.0293 0.0024 | |
| QUAVE-QCAE | 0.9842 0.002 | 0.0010 0.0001 | 0.0126 0.0034 |
| Model | Params | DSC | mIoU | MAE |
| U-Net [51] | M | 0.74743 0.05954 | 0.65463 0.01655 | 0.06525 0.00300 |
| QUAVE-U-Net | 0.77257 0.00252 | 0.63937 0.00938 | 0.06661 0.00153 | |
| U-Net++ [52] | M | 0.77960 0.00463 | 0.63790 0.01473 | 0.06635 0.00155 |
| QUAVE-U-Net++ | 0.79813 0.01066 | 0.67290 0.01094 | 0.06142 0.00244 | |
| QU-Net[35] | M | 0.76680 0.01159 | 0.62933 0.02033 | 0.06887 0.00249 |
| QUAVE-QU-Net | 0.77280 0.01022 | 0.63820 0.01585 | 0.06632 0.00273 | |
| QU-Net++[35] | M | 0.78570 0.00776 | 0.64963 0.02230 | 0.06558 0.34956 |
| QUAVE-QU-Net++ | 0.79117 0.00196 | 0.65257 0.00386 | 0.06355 0.00029 |
| Model | FID[12] | FID | IS | DSC | S-Score | mIoU | MAE |
| TarGAN [12] | 0.0912 0.0002 | 0.7009 0.1181 | 3.1879 0.2172 | 87.53 0.6474 | 76.23 9.5988 | 0.7812 0.0090 | 0.0489 0.0007 |
| QUAVE-TarGAN | 0.0893 0.0025 | 0.6295 0.0233 | 3.3452 0.0185 | 90.53 4.3608 | 84.41 4.2848 | 0.8320 0.0676 | 0.0487 0.0004 |
| QTarGAN | 0.0881 0.0021 | 0.6513 0.0144 | 2.7842 0.2628 | 83.56 1.1758 | 76.85 1.2373 | 0.7246 0.0158 | 0.0492 0.0011 |
| QUAVE-QTarGAN | 0.0902 0.0024 | 0.6253 0.0860 | 3.0228 0.0524 | 83.02 0.7664 | 78.90 1.1821 | 0.7360 0.0096 | 0.0488 0.0005 |
VI-A Evaluation
VI-A1 Reconstruction on IXI dataset
Table I shows the average and the standard deviation over different runs of the objective metrics results for the IXI dataset. The models that employ the proposed preprocessing (QUAVE) enhance the scores with both real and quaternion-valued models. These results prove that QUAVE can boost model learning without requiring modification or manipulation of models or any kind of data augmentation. Indeed, features extracted with the QWT capture salient characteristics of the image that are crucial for proper reconstruction.
VI-A2 Segmentation on Kvasir-SEG dataset
Table II reports instead objective metrics results for the segmentation task on the Kvasir-SEG dataset with RGB images. It is worth noting that the Kvasir-SEG dataset comprises RGB images, thus the real-valued U-Net is trained with three-channel color samples. Therefore, we prove that our technique, QUAVE, increases model performance also with respect to color images. Hence, QUAVE can be generalized to grayscale and color images as well, improving the accuracy of involved networks.
| Data | Model | Wavelet | SSIM | MSE | FID[12] | FID | DSC | mIoU | MAE |
IXI |
QUAVE-CAE | DWT | 0.991 0.002 | 0.001 0.001 | - | 0.010 0.001 | - | - | - |
| QWT | 0.991 0.005 | 0.001 0.001 | - | 0.003 0.001 | - | - | - | ||
| QUAVE-QCAE | DWT | 0.949 0.015 | 0.001 0.001 | - | 0.025 0.003 | - | - | - | |
| QWT | 0.984 0.002 | 0.001 0.001 | - | 0.012 0.003 | - | - | - | ||
KVASIR |
QUAVE-U-Net | DWT | - | - | - | - | 0.775 0.002 | 0.643 0.015 | 0.066 0.001 |
| QWT | - | - | - | - | 0.772 0.002 | 0.639 0.009 | 0.066 0.001 | ||
| QUAVE-QU-Net | DWT | - | - | - | - | 0.760 0.015 | 0.624 0.028 | 0.068 0.004 | |
| QWT | - | - | - | - | 0.772 0.010 | 0.638 0.015 | 0.066 0.002 | ||
| QUAVE-U-Net++ | DWT | - | - | - | - | 0.781 0.011 | 0.649 0.016 | 0.064 0.003 | |
| QWT | - | - | - | - | 0.798 0.010 | 0.672 0.010 | 0.061 0.002 | ||
| QUAVE-QU-Net++ | DWT | - | - | - | - | 0.782 0.019 | 0.645 0.026 | 0.065 0.004 | |
| QWT | - | - | - | - | 0.791 0.001 | 0.652 0.003 | 0.063 0.002 | ||
CHAOS |
QUAVE-TarGAN | DWT | - | - | 0.092 0.002 | 0.684 0.040 | 0.888 0.201 | 0.802 0.031 | 0.048 0.001 |
| QWT | - | - | 0.089 0.002 | 0.629 0.023 | 0.905 0.436 | 0.832 0.067 | 0.048 0.001 | ||
| QUAVE-QTarGAN | DWT | - | - | 0.087 0.001 | 0.649 0.069 | 0.836 0.012 | 0.724 0.017 | 0.050 0.001 | |
| QWT | - | - | 0.090 0.002 | 0.625 0.086 | 0.830 0.076 | 0.736 0.009 | 0.048 0.001 |
VI-A3 Image modality translation on CHAOS dataset
Table III reports objective metrics results for the image modality translation task on the CHAOS dataset. As in previous experiments, we apply the proposed QUAVE to data for extraction, to both the original TarGAN and a quaternion version referred to as QTargan. We can observe that both the DWT and QUAVE improve the performance with respect to the original network. The FID[12] proposed in the original TarGAN work tends to reward the images that best translate meaningful features in medical images, and our approach QUAVE beats the original processing model. Even if we consider the classical FID, QUAVE performs best among the original work. We also show this effect on a visual comparison among the predicted translation in Fig. 5 on randomly chosen samples. We display the superiority of the translation of QUAVE using red boxes. The boxes help to indicate the artifacts in the translation committed by the other methods. To conclude, we show that the data-agnostic QUAVE can be easily generalized to multimodality datasets, improving the performance of large-scale models too.
VI-B Ablation study
We adopt the QWT for our framework, consequently, we show an ablation study instead of using the DWT in QUAVE. This is possible with only a slight change in the output size of the final layers of QUAVE. From the scores in Tab. IV, we can observe that both the DWT and QUAVE approaches have a high impact on model performance, improving scores with respect to baselines. However, the best scores are, again, achieved by the models with the QUAVE method with QWT wavelet, proving its effectiveness in most of all the different tasks.
VII Conclusion
In this paper, we have introduced QUAVE a generalizable, shift-invariant, and data-agnostic framework, for exploiting salient features in medical images. The method is based on the quaternion wavelet transform and on a novel technique for extracting informative sub-bands within the Quaternion Wavelet Transform for neural models which improves networks’ performance in a variety of tasks, while improving quaternion and hypercomplex models’ generalizability. On a thorough experimental evaluation with different models in three distinct tasks, we have demonstrated the effectiveness and the generalization ability of our approach, which is able to improve performance according to a variety of metrics in each task we consider. Moreover, we have shown the flexibility of the proposed methods that can be involved in every medical and non-medical dataset, whether with grayscale, RGB images, or comprising different modalities.